OpenStudy (anonymous):

I need help solving the length and width of a rectangle by using the Pythagorean Theorem to find it. The area is 108 square inches and it has a diagonal measure of 15 inches. I am having a lot of trouble understanding how to work this formula. Any help would be appreciated.

OpenStudy (imstuck):

You're looking for the length and the width of this?

OpenStudy (anonymous):

Yes. I am trying to figure out how to achieve this. I am unsure how to.

OpenStudy (imstuck):

Working on it now before I post any advice for you, ok?

OpenStudy (anonymous):

Okay.

OpenStudy (campbell_st):

well do it this way... let the sided of the rectangle be a and b then using pythagoras a^2 + b^2 = 15^2 now the area \[a \times b = 108\] in the area formula make a the subject so \[a = \frac{108}{b}\] substitute it into pythagoras's theorem and solve for b hope it helps

OpenStudy (campbell_st):

and remember the only values that a and b can take are positives... you end up with a quartic equation that can be solve like a quadratic

OpenStudy (anonymous):

Okay. Thank you. Going to try that.

OpenStudy (campbell_st):

when you get the quartic... its an equation that's reducible to a quadratic... so use the quadratic formula to find b